OpenStudy (anonymous):

A ball is dropped from a height of 100 ft. one second later another ball is dropped from a height of 75 ft, which one hits the ground first. Solve using derivitives

OpenStudy (anonymous):

anyone have any ideas?

OpenStudy (anonymous):

Let's refer to the ball dropped from 100 ft as ball A. The ball dropped from 75 ft as ball B. Calculate the time it takes for each ball to reach the ground. Gravity makes anything dropped travel at a velocity of 32 ft/s. t = time Ball A: (32 ft/s)t = 100 ft t = 100/32 t = 3.125 s Ball B: (32 ft/s)t = 75 ft t = 75/32 t = 2.344 s Since ball A was dropped 1 second before ball B, you subtract 1 s from 3.125. So, A = 2.125, B = 2.344. Ball A reaches the ground first.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No issues.. :)